# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ Appendix D A Non-Asymptotically Valid Robust KSD Test
The decision threshold $\gamma_{n}$ is based on the concentration bound Lemma˜18, which is a deviation bound of the McDiarmid’s type (McDiarmid et al.,, 1989). This test can be viewed as a counterpart of the robust MMD test of Sun and Zou, (2023), which is also constructed using McDiarmid’s inequality. The next result, proved at the end of this section, shows its finite-sample validity under $H^{\mathrm{C}}_{0}$ as well as its consistency under $H^{\mathrm{C}}_{1}$.
@@
Remark 24.1 (Alternative deviation bounds).
@@
where $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\big{\|}\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{Y})]\big{\|}_{\mathcal{H}_{u}}$ is a Hilbert-space norm, as shown in Lemma˜16. Thus, any deviation bound for Hilbert-space norms may be applied. Examples include another McDiarmid’s bound of Gretton et al., (2012, Theorem 7), the empirical Bernstein bounds of Wolfer and Alquier, (2022, Theorem A.1) and Martinez-Taboada and Ramdas, (2024, Corollary 1), as well as the Hilbert-space valued Hoeffding bound of Pinelis, (1994, Theorem 3.5), particularly its i.i.d. variant (Chatalic et al.,, 2022, Lemma E.1).
Remark 24.2 (Justification for McDiarmid-type bound).
We opted for the McDiarmid-type bound in Lemma˜18 because, in our setting, it is tighter than the bounds from Gretton et al., (2012), Wolfer and Alquier, (2022) and Chatalic et al., (2022). In particular, while Wolfer and Alquier, (2022) claim their empirical Bernstein bound outperforms McDiarmid’s inequality, we find that this only holds for their bound tailored to translation-invariant kernels (Wolfer and Alquier,, 2022, Theorem 3.1), but not for the non-translation-invariant version (Wolfer and Alquier,, 2022, Theorem A.1), which our setting requires since we have assumed Lemma˜2. Moreover, although the bound in Lemma˜18 is worse than the empirical Bernstein bound of Martinez-Taboada and Ramdas, (2024), we find that the difference is only marginal, and the former is considerably easier to implement.
In Figure˜15, we compare R-KSD-Dev with the bootstrap-based robust-KSD test, the robust MMD test and the standard tests under the outlier-contaminated Gaussian model in Section˜5.1. We set the uncertainty radius to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ for various values of $\epsilon_{0}$, where $\tau_{\infty}$ is estimated following Section˜4.2. As expected, R-KSD-Dev controls Type-I error against Huber’s contamination with maximal proportion $\epsilon_{0}$. However, when $\epsilon>\epsilon_{0}$, it has lower power compared with robust-KSD and the MMD-based tests. This is because R-KSD-Dev uses a deviation bound to construct the decision threshold, which is a uniform bound over all data-generating distributions and is therefore conservative. For the same reason, we expect tests constructed using the alternative deviation bounds discussed in Remark˜24.1 to suffer from similar conservativeness.
@@
$=\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q}\bigg{(}\sqrt{n}\big{(}D(\mathbb{X}_{n})-D(Q,P)\big{)}>\sqrt{n}(\gamma_{n}+\theta-D(Q,P))\bigg{)}\;.$

### v5

@@ Appendix D A Non-Asymptotically Valid Robust KSD Test
The decision threshold $\gamma_{n}$ is based on the concentration bound Lemma˜18, which is a deviation bound of the McDiarmid’s type (McDiarmid et al., 1989). This test can be viewed as a counterpart of the robust MMD test of Sun and Zou (2023), which is also constructed using McDiarmid’s inequality. The next result, proved at the end of this section, shows its finite-sample validity under $H^{\mathrm{C}}_{0}$ as well as its consistency under $H^{\mathrm{C}}_{1}$.
@@
Remark 10 (Alternative deviation bounds).
@@
where $\mathbb{S}_{P}(\mathbb{X}_{n},Q)=\big\|\mathbb{E}_{\mathbf{X}\sim Q_{n}}[u_{p}(\cdot,\mathbf{X})]-\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\cdot,\mathbf{Y})]\big\|_{\mathcal{H}_{u}}$ is a Hilbert-space norm, as shown in Lemma˜16. Thus, any deviation bound for Hilbert-space norms may be applied. Examples include another McDiarmid’s bound of Gretton et al. (2012, Theorem 7), the empirical Bernstein bounds of Wolfer and Alquier (2022, Theorem A.1) and Martinez-Taboada and Ramdas (2024, Corollary 1), as well as the Hilbert-space valued Hoeffding bound of Pinelis (1994, Theorem 3.5), particularly its i.i.d. variant (Chatalic et al., 2022, Lemma E.1).
Remark 11 (Justification for McDiarmid-type bound).
We opted for the McDiarmid-type bound in Lemma˜18 because, in our setting, it is tighter than the bounds from Gretton et al. (2012), Wolfer and Alquier (2022) and Chatalic et al. (2022). In particular, while Wolfer and Alquier (2022) claim their empirical Bernstein bound outperforms McDiarmid’s inequality, we find that this only holds for their bound tailored to translation-invariant kernels (Wolfer and Alquier, 2022, Theorem 3.1), but not for the non-translation-invariant version (Wolfer and Alquier, 2022, Theorem A.1), which our setting requires since we have assumed Lemma˜2. Moreover, although the bound in Lemma˜18 is worse than the empirical Bernstein bound of Martinez-Taboada and Ramdas (2024), we find that the difference is only marginal, and the former is considerably easier to implement.
In Figure˜15, we compare R-KSD-Dev with the bootstrap-based robust-KSD test, the robust MMD test and the standard tests under the outlier-contaminated Gaussian model in Section˜5.1. We set the uncertainty radius to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ for various values of $\epsilon_{0}$, where $\tau_{\infty}$ is estimated following Section˜4.2. As expected, R-KSD-Dev controls Type-I error against Huber’s contamination with maximal proportion $\epsilon_{0}$. However, when $\epsilon>\epsilon_{0}$, it has lower power compared with robust-KSD and the MMD-based tests. This is because R-KSD-Dev uses a deviation bound to construct the decision threshold, which is a uniform bound over all data-generating distributions and is therefore conservative. For the same reason, we expect tests constructed using the alternative deviation bounds discussed in Remark˜10 to suffer from similar conservativeness.
@@
$=\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q}\Big(\sqrt{n}\big(D(\mathbb{X}_{n})-D(Q,P)\big)>\sqrt{n}\big(\gamma_{n}+\theta-D(Q,P)\big)\Big)\;.$